# Action Reasoning Models that can Reason in Space

### v1

@@ Abstract
Reasoning is central to purposeful action, yet most robotic foundation models map perception and instructions directly to control, which limits adaptability, generalization, and semantic grounding. We introduce Action Reasoning Models (ARMs), a class of vision–language–action models that integrate perception, planning, and control through a structured three-stage pipeline. Our model, MolmoAct, encodes observations and instructions into depth-aware perception tokens, generates mid-level spatial plans as editable trajectory traces, and predicts precise low-level actions, enabling explainable and steerable behavior. MolmoAct-7B-D achieves strong performance across simulation and real-world settings: $70.5\%$ zero-shot accuracy on SimplerEnv Visual Matching tasks, surpassing closed-source $\pi_{0}$ and GR00T N1; $86.6\%$ average success on LIBERO, including a $6.3\%$ gain over ThinkAct on long-horizon tasks; and in real-world fine-tuning, $10\%$ (single-arm) and $22.7\%$ (bimanual) task progression over $\pi_{0}$-FAST. It also outperforms baselines by $23.3\%$ on out-of-distribution generalization and achieves top human-preference scores for open-ended instruction following and trajectory steering. Furthermore, we release, for the first time, the MolmoAct Dataset —a mid-training robot dataset comprising over 10,000 high-quality robot trajectories across diverse scenarios and tasks. Training with this dataset yields an average 5.5% improvement in general performance over the base model. We release all model weights, training code, our collected dataset and our action reasoning dataset, establishing MolmoAct as both a state-of-the-art robotics foundation model and an open blueprint for building ARMs that transform perception into purposeful action through structured reasoning.
@@ 1 Introduction
— Barbara Tversky, Emerita Professor of Psychology at Stanford Tversky (2025)
Reasoning allows us to act with intention. Before reaching for a cup or moving through a room, we subconsciously weigh context, goals, and constraints—transforming perception into purpose. This process, grounded in our physical experience of the world, makes our actions coherent, adaptable, and explainable. For robots to operate with the same fluency, they must do more than map images and instructions to control commands. They must learn to reason.
In contrast to the rapid generalization gains seen in large language and vision models, progress in robotics has lagged behind (Duan et al., 2022; Xu et al., 2024b; Firoozi et al., 2025). Vision-Language-Action (VLA) models aim to bring similar capabilities to physical agents, but have yet to reach the same level of flexibility or robustness. Despite massive efforts in dataset collection and model scaling, today’s VLAs remain brittle and opaque—struggling to transfer across tasks, scenes, or embodiments, and offering little insight into why a robot chose one action over another (Liu et al., 2025; Pumacay et al., 2024).
@@
We introduce MolmoAct (Multimodal Open Language Model for Action), a family of completely open Action Reasoning Models (ARM) that integrate perception, planning, and control through a structured reasoning pipeline. MolmoAct learns to interpret language instructions, sense its environment, generate spatial plans, and execute them as smooth, goal-directed trajectories. The model first encodes observations and instructions into structured 2.5D representations via autoregressive prediction of depth-aware perception tokens. These tokens condition the generation of mid-level planning representations—visualized as trajectory lines in image space—which then guide the prediction of precise, low-level robot actions. This three-stage reasoning architecture enables MolmoAct to produce explainable and steerable behavior.
@@
MolmoAct is fully open in every aspect: we release the model weights, training code, and all components of our multimodal reasoning dataset. We aim for MolmoAct to be more than a high performing robotics foundation model that serves as a blueprint for building agents that reason, transforming perception into purposeful action through reasoning.
@@ 2 MolmoAct
MolmoAct is a fully open-source action reasoning model (ARM) for robotic manipulation. It builds on Molmo (Deitke et al., 2024), reusing its vision–language backbone composed of a vision encoder, a vision–language connector, and a large language model (LLM). While Molmo supports chain-of-thought reasoning for language and vision, MolmoAct extends this capability to generate structured, temporally grounded action sequences. The model is trained on a subset of the Open X-Embodiment (OXE) (O’Neill et al., 2024) mixture consisting of BC-Z (Jang et al., 2022), BridgeData V2 (Walke et al., 2023), and RT-1 Brohan et al. (2022), and is augmented with a newly collected MolmoAct Dataset.
In the following sections, we describe the VLM preliminaries (subsection 2.1), our method for tokenizing actions for vision–language actions (VLA) (subsection 2.2), the transformation of VLAs into an action reasoning model (ARM) (subsection 2.3), and our approach to action steerability via visual reasoning traces (subsection 2.4).
@@ 2.1 Vision Language Model
Our work builds on Molmo, the Multimodal Open Language Model, which follows this standard design. It employs a Vision Transformer (ViT) visual encoder, a two-layer MLP connector for projecting vision features into the language embedding space, and a decoder-only LLM backbone. In our implementation, we use vision encoders such as OpenAI’s ViT-L/14 336px CLIP (Radford et al., 2021) and ViT-SO400M/14 384px SigLIP2 (Tschannen et al., 2025), paired with open LLMs including OLMo2-7B (OLMo et al., 2024) and Qwen2.5-7B (Qwen et al., 2025). We trained MolmoAct-7B-O with a VLM backbone based on OpenCLIP and OLMo2-7B, and MolmoAct-7B-D with a backbone based on SigLIP2 and Qwen2.5-7B. For full details of our model architecture and implementation, please refer to Appendix A.
The degree of openness in the backbone components varies. SigLIP2 and Qwen2.5 do not disclose the full details of their pre-training, and are presumed to use large-scale Internet-sourced multimodal data. In contrast, OLMo2 provides open training datasets (e.g., LAION-2B/5B (Schuhmann et al., 2022), Dolma (Soldaini et al., 2024)), model weights, and complete training code. Although OpenAI’s CLIP also uses closed training data, it can be reproduced from scratch, as shown by MetaCLIP (Xu et al., 2024a). We use the model from OpenAI because it was trained for higher resolution images, and also following the previous choice from Molmo (Deitke et al., 2024). In MolmoAct, we initialize vision and language components from these open checkpoints whenever available, and use the pre-training procedure from Molmo (Deitke et al., 2024) to train the VLM on dense captioning data. Then, after vision-language alignment, we start to fine-tune MolmoAct with a subset of Open X-Embodiment (OXE) mixture and the MolmoAct Dataset. This enables full reproducibility and supports community-driven re-training and ablation studies on data curation and scaling.
@@ 2.2 Vision-Language-Action Model
Prior work represents the 256 discretized action bins with 256 distinct language tokens taken from the tail of the vocabulary. Continuous action bins, however, possess ordinal structure and local correlation, whereas arbitrary language tokens are effectively unrelated. This mismatch yields a poor initialization for learning an action codebook. We adopt a simple, order-preserving alternative that better reflects the geometry of the action space. We first identify the final 256 tokens in the Qwen2 tokenizer and, for each, use its underlying byte-level BPE symbol. These multi-character tokens admit a natural lexicographic ordering and are already sorted in order. We then assign them monotonically to the 256 bins so that adjacent bins map to adjacent symbols, and this becomes our action token vocabulary $V_{\mathrm{action}}$. This proximity-preserving initialization offers a smoother starting point for optimization and, in practice, substantially reduces training time. Retrospectively, compared to GR00T N1 ’s (NVIDIA et al., 2025) training requirement of 50,000 GPU hours, MolmoAct achieves pre-training with only 9,216 GPU hours which uses 5.4 times less compute than GR00T N1.
@@ 2.3 Action Reasoning Model
Chain-of-Thought (CoT) (Wei et al. (2022)) has been shown to significantly improve Language Models’ performance on complex task. Likewise, Multimodal Language Models (MLLM) also benefit from multimodal Chain of Thought (MCoT) (Wang et al. (2025)) in processing long multimodal contexts. However, this "think-before-you-act" paradigm is rarely present in robotic control policies. While some work attempts to incorporate reasoning to VLAs, they often emphasize high-level plans or textual hints (Sun et al., 2024; Zawalski et al., 2024) (for example, subtask or direction signals) and object-centric cues (for example, detections or boxes) (Li et al., 2025; Gu et al., 2023). While useful, they ignore two crucial aspects for precise control: depth perception and precise motion plans. First, most VLM backbones are trained on RGB imagery and receive no direct supervision about depth, yet accurate 3D understanding is central to physical interaction. Second, linguistic directives are inherently ambiguous with respect to magnitudes, scales, and endpoints; the same phrase can map to many distinct trajectories. Our premise is therefore to make the model reason explicitly about depth and about a concrete, image-aligned motion sketch before emitting low-level actions.
In MolmoAct, we incorporate intermediate reasoning not through pure language-based reasoning; rather, we teach our models to reason in space. We cast visuomotor control as autoregressive sequence modeling augmented with action CoT as to achieve reasoning in space. Conditioned on images and instructions, the autogregressive action reasoning model first generates a short sequence of depth perception tokens that summarize depth perception of the current observation, followed by a compact visual reasoning trace of the intended end-effector motion, and only then produces the action tokens for the next control step. This approach mirrors how textual CoT is performed; however, with action CoT, every intermediate step is explicitly grounded in space.
@@
Depth estimation is a key aspect for spatial understanding and robot action prediction. Most traditional VLMs and VLAs only condition their outputs solely on the RGB image input and text instruction, which worsens the prediction for tasks that heavily require spatial understanding. Prior work (Bigverdi et al., 2024) has shown that depth perception tokens are effective in enhancing chain-of-thought reasoning for visual–spatial tasks. Building on this insight, we incorporate depth token prediction as a key component of our action reasoning in space, enabling more accurate control of the robot in 3D environments and helping to bridge this critical gap. We now examine how each intermediate step in the reasoning chain is formulated:
@@
$V_{\mathrm{depth}}=\{\langle\mathrm{DEPTH\_START}\rangle,\langle\mathrm{DEPTH\_END}\rangle\}\;\cup\;\{\langle\mathrm{DEPTH}\_k\rangle\}_{k=1}^{N}$
@@
$\mathbf{d}=\big{(}\langle\mathrm{DEPTH\_START}\rangle\;\langle\mathrm{DEPTH}\_{z^{\mathrm{depth}}_{1}}\rangle\;\ldots\;\langle\mathrm{DEPTH}\_{z^{\mathrm{depth}}_{M}}\rangle\;\langle\mathrm{DEPTH\_END}\rangle\big{)}\in V_{\mathrm{depth}}.$
@@
Planning in advance is often crucial for robots to complete tasks. Instead of planning through text description, we choose to plan via the trajectory of the robot’s end effector projected into the image plane. This trajectory is a more grounded and precise way than language to show how the robot should move from the current observation until task completion.
@@
$\boldsymbol{\tau}\;=\;(p_{1},\ldots,p_{L}),\qquad p_{i}=(u_{i},v_{i}).$
@@
$p(\mathbf{d},\boldsymbol{\tau},\mathbf{a}\,\big{\lvert}\,I,T)=\>\prod_{i=1}^{M+2}p\bigl{(}d_{i}\mid I,T,\mathbf{d}_{<i}\bigr{)}\;\times\;\prod_{j=1}^{L}p\bigl{(}\tau_{j}\mid I,T,\mathbf{d},\boldsymbol{\tau}_{<j}\bigr{)}\;\times\;\prod_{k=1}^{D}p\bigl{(}a_{k}\mid I,T,\mathbf{d},\boldsymbol{\tau},\mathbf{a}_{<k}\bigr{)}.$
@@ 2.4 Action Steerability via Visual Reasoning Trace
$p(\mathbf{a}\mid I^{+},T)=\;\prod_{k=1}^{D}p\bigl{(}a_{k}\mid I^{+},T,\mathbf{a}_{<k}\bigr{)}.$
@@ 3 Data Curation and Generation
MolmoAct is trained on a diverse set of datasets. During Pre-training, MolmoAct is trained on Multimodal Web data, Auxiliary Robot data as well as Action Reasoning data. Furthermore, we collected and trained wtih the MolmoAct Dataset for the Mid-training stage. Below, we describe each dataset and its collection process; further details and examples are provided in the Appendix E.
@@ 4 Training Recipe
MolmoAct is first pre-trained on action reasoning data curated from a subset of the OXE dataset, along with the auxiliary robot dataset and multimodal web dataset. To further enhance its capabilities, we mid-train the model on the MolmoAct Dataset before post-training it for specific downstream tasks and embodiments. In this section, we describe the different data mixtures used at each stage of MolmoAct ’s training.
@@ 4.1 Pre-training
In the first training stage, MolmoAct is pre-trained on a mixture of action reasoning data, auxiliary robot data, and multimodal web data. For robot data, we use a subset of OXE comprising RT-1, BridgeData V2, and BC-Z, totaling 10.5M samples, which we convert into action reasoning data using our action CoT formulation. We also include Auxiliary Robot Data—atomic depth data (1.5M), atomic trace data (1.5M), and trajectory-conditioned action data (10.5M), and co-train with 2M samples of multimodal web data. During pre-training, data is sampled at the following rates: RT-1 (20%), BridgeData V2 (12.5%), BC-Z (7.5%) for both action reasoning and trajectory-conditioned datasets, 7.5% from the auxiliary depth and visual trace datasets, and 5% from multimodal web data as shown in Figure 3. The whole data mixture used for pre-training MolmoAct totals up to 26.3M samples.
To pre-train MolmoAct with the data mixture mentioned above, we use 256 H100s to train the model with 100k gradient steps using a batch size of 512 takes up 9,728 GPU hours. At each training step, a batch of data pairs is extracted at random timesteps from the shuffled robot demonstrations or randomly from the multimodal web data. The data pairs are drawn based on the sampling rate defined above.

### v2

@@ Abstract
Reasoning is central to purposeful action, yet most robotic foundation models map perception and instructions directly to control, which limits adaptability, generalization, and semantic grounding. We introduce Action Reasoning Models (ARMs), a class of robotic foundation models that integrates perception, planning, and control through a structured three-stage pipeline. Our model, MolmoAct, encodes observations and instructions into depth-aware perception tokens, generates mid-level spatial plans as editable trajectory traces, and predicts precise low-level actions, enabling explainable and steerable behavior. MolmoAct-7B-D achieves strong performance across simulation and real-world settings: $70.5\%$ zero-shot accuracy on SimplerEnv Visual Matching tasks, surpassing closed-source $\pi_{0}$ and GR00T N1; $86.6\%$ average success on LIBERO, including a $6.3\%$ gain over ThinkAct on long-horizon tasks; and in real-world fine-tuning, $10\%$ (single-arm) and $22.7\%$ (bimanual) task progression over $\pi_{0}$-FAST. It also outperforms baselines by $23.3\%$ on out-of-distribution generalization and achieves top human-preference scores for open-ended instruction following and trajectory steering. Furthermore, we release, for the first time, the MolmoAct Dataset —a mid-training robot dataset comprising over 10,000 high-quality robot trajectories across diverse scenarios and tasks. Training with this dataset yields an average 5.5% improvement in general performance over the base model. We release all model weights, training code, MolmoAct Dataset and our action reasoning dataset, establishing MolmoAct as both a state-of-the-art robotics foundation model and an open blueprint for building ARMs that transform perception into purposeful action through grounded reasoning.
@@ 1 Introduction
— Barbara Tversky, Emerita Professor of Psychology at Stanford (Tversky, 2025)
Reasoning allows us to act with intention. Before reaching for a cup or moving through a room, we subconsciously weigh context, goals, and constraints—transforming perception into purpose. This process, grounded in our physical experience of the world, makes our actions coherent, adaptable, and explainable. For robots to operate with the same fluency, they must do more than map images and instructions to robot control. They must learn to reason.
In contrast to the rapid generalization gains seen in large language and vision models, progress in robotics has lagged behind (Duan et al., 2022; Xu et al., 2024b; Firoozi et al., 2025). Vision-Language-Action (VLA) models (Black et al., ; Kim et al., 2024; Team et al., 2025; NVIDIA et al., 2025; Yang et al., 2025b) aim to bring similar capabilities to physical agents, but have yet to reach the same level of flexibility or robustness. Despite massive efforts in dataset collection and model scaling, today’s VLAs remain brittle and opaque—struggling to transfer across tasks, scenes, or embodiments, and offering little insight into why a robot chose one action over another (Liu et al., 2025; Pumacay et al., 2024).
@@
We introduce MolmoAct (Multimodal Open Language Model for Action), a family of completely open Action Reasoning Models (ARM) that integrate perception, planning, and control through a structured reasoning pipeline. MolmoAct learns to interpret language instructions, sense its environment, generate spatial plans, and execute them as smooth, goal-directed trajectories. The model first encodes observations and instructions into structured 2.5D representations via autoregressive prediction of depth-aware perception tokens. These tokens condition the generation of mid-level planning representations, which, when visualized as visual traces in image space, guide the prediction of precise, low-level robot actions. This three-stage reasoning architecture enables MolmoAct to produce explainable and steerable behavior as shown in Figure 1.
@@
MolmoAct is fully open in every aspect: we release the model weights, training code, and all components of our action reasoning dataset. We aim for MolmoAct to be more than a high-performing robotics foundation model that serves as a blueprint for building agents that reason, transforming perception into purposeful action through reasoning.
@@ 2 MolmoAct
MolmoAct is a fully open-source action reasoning model (ARM) for robotic manipulation. It builds on Molmo (Deitke et al., 2024), reusing its vision–language backbone composed of a vision encoder, a vision–language connector, and a large language model (LLM). While Molmo supports chain-of-thought reasoning for language and vision, MolmoAct extends this capability to generate grounded action sequences. The model is trained on a subset of the Open X-Embodiment (OXE) (O’Neill et al., 2024) mixture consisting of BC-Z (Jang et al., 2022), BridgeData V2 (Walke et al., 2023), and RT-1 (Brohan et al., 2022), and our in-house collected MolmoAct Dataset.
In the following sections, we describe the VLM preliminaries (subsection 2.1), our method to adapt VLMs for action prediction via action tokenization(subsection 2.2), how we transform Molmo into an action reasoning model (ARM, subsection 2.3), and our approach to steer action by visual reasoning traces (subsection 2.4).
@@ 2.1 Vision Language Model
Our work builds on Molmo, the Multimodal Open Language Model, which follows this standard design. It employs a Vision Transformer (ViT) visual encoder, a two-layer MLP connector for projecting vision features into the language embedding space, and a decoder-only LLM backbone. In our implementation, we use vision encoders such as OpenAI ViT-L/14 336px CLIP (Radford et al., 2021) and ViT-SO400M/14 384px SigLIP2 (Tschannen et al., 2025), paired with open LLMs including OLMo2-7B (OLMo et al., 2024) and Qwen2.5-7B (Qwen et al., 2025). We trained MolmoAct-7B-O with a VLM backbone based on OpenCLIP and OLMo2-7B, and MolmoAct-7B-D with a backbone based on SigLIP2 and Qwen2.5-7B. For full details of our model architecture and implementation, please refer to Appendix A.
The degree of openness in the backbone components varies. SigLIP2 and Qwen2.5 do not disclose the full details of their pre-training, and are presumed to use large-scale Internet-sourced multimodal data. In contrast, OLMo2 provides open training datasets (e.g., LAION-2B/5B (Schuhmann et al., 2022), Dolma (Soldaini et al., 2024)), model weights, and complete training code. Although OpenAI CLIP also uses closed training data, it can be reproduced from scratch, as shown by MetaCLIP (Xu et al., 2024a). We use the model from OpenAI because it was trained for higher resolution images, and also following the previous choice from Molmo (Deitke et al., 2024). In MolmoAct, we initialize vision and language components from these open checkpoints whenever available, and use the pre-training procedure from Molmo (Deitke et al., 2024) to train the VLM on dense captioning data. Then, after vision-language alignment, we start to fine-tune MolmoAct with a subset of Open X-Embodiment (OXE) mixture and the MolmoAct Dataset. This enables full reproducibility and supports community-driven re-training and ablation studies on data curation and scaling.
@@ 2.2 Vision-Language-Action Model
Prior work represents the 256 discretized action bins with 256 distinct language tokens taken from the tail of the vocabulary. Continuous action bins, however, possess ordinal structure and local correlation, whereas arbitrary language tokens are effectively unrelated. This mismatch yields a poor initialization for learning an action codebook. We adopt a simple alternative that better reflects the geometry of the action space. We first identify the final 256 tokens in the Qwen2 tokenizer and, for each, use its underlying byte-level BPE symbol. We then assign them monotonically to the 256 bins so that adjacent bins map to adjacent symbols, and this becomes our action token vocabulary $V_{\mathrm{action}}$. We notice that these BPE symbols have a better initialization for action token embeddings by sharing similar characters between adjacent action bins. This similarity-preserving initialization offers a smoother starting point for optimization and, in practice, substantially reduces training time. In contrast to GR00T N1’s (NVIDIA et al., 2025) training time of 50,000 GPU hours, MolmoAct achieves pre-training with only 9,216 GPU hours: over a 5x reduction.
@@ 2.3 Action Reasoning Model
Chain-of-Thought (CoT) (Wei et al., 2022) has been shown to significantly improve Language Models’ performance on complex tasks. Likewise, Multimodal Language Models (MLLM) also benefit from multimodal Chain of Thought (MCoT) (Lai and Nissim, 2024) in processing long multimodal contexts. However, this "think-before-you-act" paradigm is rarely present in robotic control policies. While some work attempts to incorporate reasoning to VLAs, they focus on high-level language reasoning (Sun et al., 2024; Zawalski et al., 2024; Intelligence et al., 2025) such as decomposing a high-level semantic task into subtasks. While useful, they ignore two crucial aspects for precise control: depth perception and precise motion planning. First, most VLMs are trained solely on RGB images and hence lack the ability of depth estimation and 3D understanding, which is critical for robotic manipulation. Moreover, attempting to distill complex 3D trajectories into linguistic descriptions often results in significant loss of spatial and temporal information.
Contrast to previous approaches, MolmoAct does not incorporate intermediate reasoning through language; rather, we teach our models to reason in space. Conditioned on images and instructions, the model autogregressively generates a sequence of depth perception tokens, followed by visual reasoning trace of the intended end-effector motion, before predicting the action tokens.
@@
Depth estimation is a key aspect for spatial understanding and robot action prediction. Most conventional VLMs and VLAs condition their outputs solely on the RGB image input and text instruction. Lacking depth estimation, they fail on tasks that require spatial understanding. Prior work (Bigverdi et al., 2024) has shown that depth perception tokens are effective in enhancing chain-of-thought reasoning for visual–spatial tasks. Building on this insight, we leverage depth estimation as a key component to enable fine-grained robotics control in 3D environments. We now examine how each intermediate step in the reasoning chain is formulated:
@@
$V_{\mathrm{depth}}=\{\langle\mathrm{DEPTH\_START}\rangle,\langle\mathrm{DEPTH\_END}\rangle\}\;\cup\;\{\langle\mathrm{DEPTH}\_k\rangle\}_{k=1}^{N}$ (1)
@@
$\mathbf{d}=\big{(}\langle\mathrm{DEPTH\_START}\rangle\;\langle\mathrm{DEPTH}\_{z^{\mathrm{depth}}_{1}}\rangle\;\ldots\;\langle\mathrm{DEPTH}\_{z^{\mathrm{depth}}_{M}}\rangle\;\langle\mathrm{DEPTH\_END}\rangle\big{)}\in V_{\mathrm{depth}}.$ (2)
@@
Planning is a crucial component in robotics. Instead of planning through subtask decomposition in language, we predict intermediate two-dimensional representation to better align visual inputs and control outputs across diverse robots and tasks. In particular, we generate the 2D trajectory of the end-effector and train the model to predict this trajectory alongside the next action command, a strategy shown to be effective in prior work (Li et al., 2025; Zheng et al., 2024; Gu et al., 2023; Niu et al., 2024). These predicted waypoints help align each action to precise end-effector locations, enabling the model to focus on fine-grained localization and thereby improving action-prediction accuracy. To achieve this, we introduce Visual Reasoning Trace.
@@
$\boldsymbol{\tau}\;=\;(p_{1},\ldots,p_{L}),\qquad p_{i}=(u_{i},v_{i}).$ (3)
@@
$p(\mathbf{d},\boldsymbol{\tau},\mathbf{a}\,\big{\lvert}\,I,T)=\>\prod_{i=1}^{M+2}p\bigl{(}d_{i}\mid I,T,\mathbf{d}_{<i}\bigr{)}\;\times\;\prod_{j=1}^{L}p\bigl{(}\tau_{j}\mid I,T,\mathbf{d},\boldsymbol{\tau}_{<j}\bigr{)}\;\times\;\prod_{k=1}^{D}p\bigl{(}a_{k}\mid I,T,\mathbf{d},\boldsymbol{\tau},\mathbf{a}_{<k}\bigr{)}.$ (4)
@@ 2.4 Action Steerability via Visual Reasoning Trace
$p(\mathbf{a}\mid I^{+},T)=\;\prod_{k=1}^{D}p\bigl{(}a_{k}\mid I^{+},T,\mathbf{a}_{<k}\bigr{)}.$ (5)
@@ 3 Data Curation and Generation
MolmoAct is trained on a diverse set of datasets. During Pre-training, MolmoAct is trained on Multimodal Web data, Auxiliary Robot Data as well as Action Reasoning Data. Furthermore, we collected and trained with the MolmoAct Dataset for the Mid-training stage. Below, we describe each dataset and its collection process; further details and examples are provided in the Appendix E.
@@ 4 Training Recipe
MolmoAct is first pre-trained on action reasoning data curated from a subset of the OXE dataset, along with the auxiliary robot data and multimodal web data. To further enhance its capabilities, we mid-train the model on the MolmoAct Dataset before post-training it for specific downstream tasks and embodiments. In this section, we describe the different data mixtures and training configurations used at each stage of MolmoAct ’s training (as shown in Figure 2.
@@ 4.1 Pre-training
In the first training stage, MolmoAct is pre-trained on a mixture of action reasoning data, auxiliary robot data, and multimodal web data. For all robot data, we use a subset of OXE comprising RT-1, BridgeData V2, and BC-Z, totaling 10.5M samples, which we convert into action reasoning data using our reasoning in space formulation. We also include auxiliary robot data—auxiliary depth data (1.5M), auxiliary trace data (1.5M), and trajectory-conditioned action data (10.5M), and co-train with 2M samples of multimodal web data. During pre-training, data is sampled at the following rates: RT-1 (20%), BridgeData V2 (12.5%), BC-Z (7.5%) for both action reasoning and trajectory-conditioned data, 7.5% from the auxiliary depth and trace data, and 5% from multimodal web data as shown in Figure 3. The whole data mixture used for pre-training MolmoAct totals up to 26.3M samples.
To pre-train MolmoAct with the data mixture mentioned above, we use 256 H100s to train the model with 100k gradient steps using a batch size of 512, which takes around 9,728 GPU hours. At each training step, a batch of data pairs is drawn randomly from the entire data mixture by their assigned sampling rate defined above. Hyperparameter details are listed in Appendix B.